OpenStudy (anonymous):

Use the quadratic formula to solve the equation? 4x2 – 6x – 10 = 0 is anybody able to show me work?

OpenStudy (anonymous):

the quadratic formula is -b (±)Sqrt[b-4ac]/2a\[-b \pm \frac{ \sqrt{b-4ac}}{ 2a}\]

OpenStudy (anonymous):

where ax^2 +bx +c just substitute and you will get your answer.

jimthompson5910 (jim_thompson5910):

\[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(-6)\pm\sqrt{(-6)^2-4(4)(-10)}}{2(4)}\] \[\Large x = \frac{6\pm\sqrt{36-(-160)}}{8}\] I'll let you finish.

OpenStudy (anonymous):

I have all that work! I just get confused with the last part? like would i subtract 36-(-160)

OpenStudy (anonymous):

36--160? well double negative is positive, so? 36+160

jimthompson5910 (jim_thompson5910):

36-(-160) is the same as 36+160

OpenStudy (anonymous):

so it would become 196?

jimthompson5910 (jim_thompson5910):

it would

OpenStudy (anonymous):